# Structure-preserving space discretization of differential and nonlocal constitutive relations for port-Hamiltonian systems

### v1

@@ 4.1.2 Finite Element approximation
Let us now discretize (40). This system is composed of two state variables $\sigma$ and $v$. Due to the choice of causality, the spatial derivative of $v$ is evaluated strongly, hence has to be discretized in $H^{1}$. Moreover, due to operator in front of $\sigma$, it has to be discretized in $H^{1}$ as well. Let us choose $(\phi_{i})_{i\in[1,N]}$ the set of $P^{1}$ Finite Element family over the points $a=x_{1}<x_{2}<...<x_{N}=b$ satisfying $\forall i,j\in[1,N],\,\phi_{i}(x_{j})=\delta_{i}(j)$ and write the discretized versions of $\sigma$ and $v$ as follows:
@@
$\sum_{j=1}^{N}\frac{1}{E}\left[\int_{a}^{b}\phi_{i}\phi_{j}\,\textup{d}x+\ell^{2}\,\int_{a}^{b}\partial_{x}\phi_{i}\,\partial_{x}\phi_{j}\,\textup{d}x\right]\frac{\textup{d}}{\textup{d}t}\overline{\sigma}_{j}=&\sum_{j=1}^{N}\int_{a}^{b}\phi_{i}\,\partial_{x}\phi_{j}\,\textup{d}x\,\overline{v}_{j}\\ &+\delta_{1}(i)\frac{\ell^{2}}{E}\frac{\textup{d}}{\textup{d}t}(u_{L})_{1}+\delta_{N}(i)\frac{\ell^{2}}{E}\frac{\textup{d}}{\textup{d}t}(u_{L})_{2},\\ \sum_{j=1}^{N}\int_{a}^{b}\phi_{i}\,\rho\,\phi_{j}\,\textup{d}x\,\frac{\textup{d}}{\textup{d}t}\overline{v}_{j}=&-\sum_{j=1}^{N}\int_{a}^{b}\partial_{x}\phi_{i}\,\phi_{j}\,\textup{d}x\,\overline{\sigma}_{j}\\ &+\delta_{1}(i)\,(u_{D})_{1}+\delta_{N}(i)\,(u_{D})_{2}.\end{aligned}\right.$
@@
$\frac{1}{E}(\bm{M}+\ell^{2}\,\bm{K})\frac{\textup{d}}{\textup{d}t}\overline{\sigma}&=\bm{D}\,\overline{v}+\frac{\ell^{2}}{E}\bm{B}\,\frac{\textup{d}}{\textup{d}t}u_{L},\\ \bm{M}^{\rho}\frac{\textup{d}}{\textup{d}t}\overline{v}&=-\bm{D}^{\top}\overline{\sigma}+\bm{B}\,u_{D}.\end{aligned}\right.$
@@
$\frac{1}{E}(\bm{M}+\ell^{2}\,\bm{K}+\ell\,\bm{BB}^{\top})\frac{\textup{d}}{\textup{d}t}\overline{\sigma}&=D\,\overline{v},\\ (\bm{M}^{\rho}+\ell\,\bm{B}\bm{M}^{\partial}_{\rho}\bm{B}^{\top})\frac{\textup{d}}{\textup{d}t}\overline{v}&=-D^{\top}\overline{\sigma}.\end{aligned}\right.$
@@ 4.3 Incompressible Navier–Stokes equations
Then, such state satisfies the following power balance:
@@ 4.3.1 Weak formulation
$-\rho_{0}\int_{\Omega}\phi_{1}\Delta\partial_{t}\psi\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\phi_{1}\operatorname{div}(\omega\,\,\textbf{{grad}}^{\perp}(\psi))\,\textup{d}x-\mu\int_{\Omega}\phi_{1}\,\Delta^{2}\psi\,\textup{d}x,\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\phi_{2}\operatorname{div}(\textbf{{grad}}^{\perp}(\psi)\,\,\omega)\,\textup{d}x+\mu\int_{\Omega}\phi_{2}\,\Delta\omega\,\textup{d}x.\end{aligned}\right.$ (45)
@@
$\rho_{0}\int_{\Omega}\textbf{{grad}}(\phi_{1})\cdot\partial_{t}\textbf{{grad}}(\psi)\,\textup{d}x&={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\partial_{t}\,\bm{n}\cdot\textbf{{grad}}(\psi)\,\textup{d}s}+\int_{\Omega}\rho_{0}\,\omega\,\,\textbf{{grad}}(\phi_{1})\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}x\\ &\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\,\omega\,\bm{n}\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}s}-\mu\int_{\Omega}\Delta\phi_{1}\,\Delta\psi\,\textup{d}x\\ &\quad{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\mu\int_{\partial\Omega}\phi_{1}\,\bm{n}\cdot\textbf{{grad}}(\Delta\psi)\,\textup{d}s+\mu\int_{\partial\Omega}\Delta\psi\,\bm{n}\cdot\textbf{{grad}}(\phi_{1})\,\textup{d}s},\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\textbf{{grad}}^{\perp}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\,\psi\,\textup{d}x{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\,-\int_{\partial\Omega}\rho_{0}\,\phi_{2}\,\psi\,\bm{t}\cdot\textbf{{grad}}(\omega)\,\textup{d}s}\\ &\quad-\mu\int_{\Omega}\textbf{{grad}}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\textup{d}x{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}+\mu\int_{\partial\Omega}\phi_{2}\,\bm{n}\cdot\textbf{{grad}}(\omega)\,\textup{d}s}.\end{aligned}\right.$ (46)
@@
$\rho_{0}\int_{\Omega}\textbf{{grad}}(\phi_{1})\cdot\partial_{t}\textbf{{grad}}(\psi)\,\textup{d}x&={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{0}\,\int_{\partial\Omega}\phi_{1}\,\partial_{t}\,u_{L}\,\textup{d}s}+\int_{\Omega}\rho_{0}\,\omega\,\,\textbf{{grad}}(\phi_{1})\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}x\\ &\quad+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\,\omega\,u_{D}^{1}\,\textup{d}s}-\mu\int_{\Omega}\Delta\phi_{1}\,\Delta\psi\,\textup{d}x\\ &\quad{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\mu\int_{\partial\Omega}\phi_{1}\,u_{D}^{2}\,\textup{d}s+\mu\int_{\partial\Omega}u_{D}^{3}\,\bm{n}\cdot\textbf{{grad}}(\phi_{1})\,\textup{d}s},\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\textbf{{grad}}^{\perp}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\,\psi\,\textup{d}x{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+\int_{\partial\Omega}\rho_{0}\,\phi_{2}\,\psi\,u_{D}^{4}\textup{d}s}\\ &\quad-\mu\int_{\Omega}\textbf{{grad}}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\textup{d}x{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}+\mu\int_{\partial\Omega}\phi_{2}\,u_{D}^{5}\,\textup{d}s}.\end{aligned}\right.$ (47)
@@ 4.3.2 Finite Element approximation
Due to the dissipation term $\int_{\Omega}\Delta\psi\Delta\phi_{1}\,\textup{d}x$, $\psi$ has to be in $H^{2}$, hence let us choose an Argyris finite element family $(\phi^{1}_{i})_{i\in[1,N_{\psi}]}$. Given that Argyris elements are polynomials of degree 5 and the relation $-\Delta\psi=\omega$, let us choose $H^{1}$ conforming $P^{3}$ Lagrange finite element family $(\phi^{2}_{i})_{i\in[1,N_{\omega}]}$ over $\Omega$. The discretized state variables are denoted by $\hat{\omega},\hat{\psi}$ and are defined as: $\hat{\psi}(t,x)=\sum_{i=1}^{N_{\psi}}\phi_{i}^{1}(x)\overline{\psi}_{i}(t)$, $\hat{\omega}(t,x)=\sum_{i=1}^{N_{\omega}}\phi^{2}_{i}(x)\overline{\omega}_{i}(t)$.
@@
$\rho_{0}\,\frac{\textup{d}}{\textup{d}t}\,\begin{bmatrix}\bm{K}&0\\ 0&\bm{M}\end{bmatrix}\begin{pmatrix}\overline{\psi}\\ \overline{\omega}\end{pmatrix}=&\begin{bmatrix}\begin{bmatrix}\bm{D}^{1}(\overline{\omega})&0\\ 0&-\bm{D}^{2}(\overline{\psi})\end{bmatrix}-\mu\begin{bmatrix}\bm{R}^{1}&0\\ 0&\bm{R}^{2}\end{bmatrix}\end{bmatrix}\begin{pmatrix}\overline{\psi}\\ \overline{\omega}\end{pmatrix}\\ &+\begin{bmatrix}\mu\,\bm{B}^{1}&\mu\,\bm{B}^{3}&\bm{B}^{3}\\ 0&0&0\end{bmatrix}\begin{pmatrix}u_{D}^{1}\\ u_{D}^{5}\\ \tilde{u}_{D}\end{pmatrix}+\mu\begin{bmatrix}0\\ \bm{B}^{5}\end{bmatrix}u_{D}^{5},\\ 0=&-\bm{B}^{1\top}\overline{\psi},\qquad\text{(No slip B.C.)}\\ 0=&-\bm{B}^{3\top}\overline{\psi},\qquad\text{(Impermeable B.C.)}\\ \bm{M}^{\partial}u_{D}^{1}=&\bm{B}^{5\top}\overline{\omega},\qquad\text{(Vorticity generation)}\end{aligned}\right.$ (51)

### v2

@@ 4.1.2 Finite Element approximation
Let us now discretize (39). This system is composed of two state variables $\sigma$ and $v$. Due to the choice of causality, the spatial derivative of $v$ is evaluated strongly, hence has to be discretized in $H^{1}$. Moreover, due to the operator in front of $\sigma$, it has to be discretized in $H^{1}$ as well. Let us choose $(\phi_{i})_{i\in[1,N]}$ the set of $P^{1}$ Finite Element family over the points $a=x_{1}<x_{2}<...<x_{N}=b$ satisfying $\forall i,j\in[1,N],\,\phi_{i}(x_{j})=\delta_{i}(j)$ and write the discretized versions of $\sigma$ and $v$ as follows:
@@
$\left\{\begin{aligned} \sum_{j=1}^{N}\frac{1}{E}\left[\int_{a}^{b}\phi_{i}\phi_{j}\,\textup{d}x+\ell^{2}\,\int_{a}^{b}\partial_{x}\phi_{i}\,\partial_{x}\phi_{j}\,\textup{d}x\right]\frac{\textup{d}}{\textup{d}t}\overline{\sigma}_{j}=&\sum_{j=1}^{N}\int_{a}^{b}\phi_{i}\,\partial_{x}\phi_{j}\,\textup{d}x\,\overline{v}_{j}\\ &+\delta_{1}(i)\frac{\ell^{2}}{E}\frac{\textup{d}}{\textup{d}t}(u_{L})_{1}+\delta_{N}(i)\frac{\ell^{2}}{E}\frac{\textup{d}}{\textup{d}t}(u_{L})_{2},\\ \sum_{j=1}^{N}\int_{a}^{b}\phi_{i}\,\rho\,\phi_{j}\,\textup{d}x\,\frac{\textup{d}}{\textup{d}t}\overline{v}_{j}=&-\sum_{j=1}^{N}\int_{a}^{b}\partial_{x}\phi_{i}\,\phi_{j}\,\textup{d}x\,\overline{\sigma}_{j}\\ &+\delta_{1}(i)\,(u_{D})_{1}+\delta_{N}(i)\,(u_{D})_{2}.\end{aligned}\right$.
@@
$\left\{\begin{aligned} \frac{1}{E}(\bm{M}+\ell^{2}\,\bm{K})\frac{\textup{d}}{\textup{d}t}\overline{\sigma}&=\bm{D}\,\overline{v}+\frac{\ell^{2}}{E}\bm{B}\,\frac{\textup{d}}{\textup{d}t}u_{L},\\ \bm{M}^{\rho}\frac{\textup{d}}{\textup{d}t}\overline{v}&=-\bm{D}^{\top}\overline{\sigma}+\bm{B}\,u_{D}.\end{aligned}\right$.
@@
Remark 14.
In the setting of Section 2.1, the symmetry of $\bm{P}^{*}\bm{S}$ (where a star denotes the adjoint) is expressed with the usual Euclidean inner product on $\mathbb{R}^{n+n_{L}}$, hence $\bm{P}^{*}=\bm{P}^{\top}$, and checking the symmetry of the operator $\bm{P}^{*}\bm{S}$ amounts to checking the symmetry of the matrix $\bm{P}^{\top}\bm{S}$. However, when applying the Finite Element Method, an invertible mass matrix $\mathbb{M}$ appears on the left-hand side of (4b). The symmetry condition is then expressed using the weighted inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{-1}}$, and checking the symmetry of $\bm{P}^{*}\bm{S}$ is then achieved by checking the symmetry of the matrix $\bm{P}^{\top}\mathbb{M}^{-1}\bm{S}$.
@@
$\left\{\begin{aligned} \frac{1}{E}(\bm{M}+\ell^{2}\,\bm{K}+\ell\,\bm{BB}^{\top})\frac{\textup{d}}{\textup{d}t}\overline{\sigma}&=D\,\overline{v},\\ (\bm{M}^{\rho}+\ell\,\bm{B}\bm{M}^{\partial}_{\rho}\bm{B}^{\top})\frac{\textup{d}}{\textup{d}t}\overline{v}&=-D^{\top}\overline{\sigma}.\end{aligned}\right$.
@@ 4.2 Implicit Euler-Bernoulli beam
Proof.
First, following Lemma 3 we have $H^{d}_{1}:=\frac{1}{2}\begin{pmatrix}\overline{\sigma}\\ \overline{v}\\ y_{L}\end{pmatrix}^{\top}\begin{pmatrix}D^{-1}\bm{M}&0&0\\ 0&\rho h\bm{M}+\frac{\rho h^{3}}{12}\bm{K}&-\frac{\rho h^{3}}{12}\bm{B}\\ 0&-\frac{\rho h^{3}}{12}\bm{B}^{\top}&0&\\ \end{pmatrix}\begin{pmatrix}\overline{\sigma}\\ \overline{v}\\ y_{L}\end{pmatrix}=\frac{D^{-1}}{2}\overline{\sigma}^{\top}\bm{M}\overline{\sigma}+\frac{\rho h}{2}\left(\overline{v}^{\top}\bm{M}\overline{v}+\frac{h^{2}}{12}\overline{v}^{\top}\bm{K}\overline{v}\right)-\frac{\rho h^{3}}{12}y_{L}^{\top}\bm{B}^{\top}\overline{v}$, the power balance is then obtained by a direct computation.
Regarding $H^{d}_{2}$, we have:
$H_{2}^{d}(\overline{\sigma},\overline{v}):$ $=H(\sigma^{d},v^{d})=\frac{1}{2}\int_{a}^{b}\rho h\left((v^{d})^{2}+\frac{h^{2}}{12}\left(\frac{\partial}{\partial x}v^{d}\right)^{2}\right)+D^{-1}\,(\sigma^{d})^{2}\,\textup{d}x,$
$=\frac{1}{2}\sum_{i}\sum_{j}\int_{a}^{b}\rho h\left(\phi_{i}\phi_{j}+\frac{h^{2}}{12}\left(\frac{\partial}{\partial x}\phi_{i}\right)\left(\frac{\partial}{\partial x}\phi_{j}\right)\right)\,\overline{v}_{i}\overline{v}_{j}\,\textup{d}x$
$\quad+\sum_{i}\sum_{j}\int_{a}^{b}D^{-1}\,\phi_{i}\phi_{j}\,\,\overline{\sigma}_{i}\overline{\sigma}_{j}\,\textup{d}x,$
$=\frac{D^{-1}}{2}\overline{\sigma}^{\top}\bm{M}\overline{\sigma}+\frac{\rho h}{2}\left(\overline{v}^{\top}\bm{M}\overline{v}+\frac{h^{2}}{12}\overline{v}^{\top}\bm{K}\overline{v}\right),$
and the power balance follows by direct computation. ∎
@@ 4.3 Incompressible Navier–Stokes equations
Then, such a state satisfies the following power balance:
@@
Proof.
The result is obtained simply by using the previously stated Theorem 14 and replacing $\omega$ by the fixed state $\omega_{0}$ for the kinetic energy, and $\psi$ by the fixed state $\psi_{0}$ in the enstrophy balance. ∎
@@ 4.3.1 Weak formulation
$\left\{\begin{aligned} -\rho_{0}\int_{\Omega}\phi_{1}\Delta\partial_{t}\psi\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\phi_{1}\operatorname{div}(\omega\,\,\textbf{{grad}}^{\perp}(\psi))\,\textup{d}x-\mu\int_{\Omega}\phi_{1}\,\Delta^{2}\psi\,\textup{d}x,\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\phi_{2}\operatorname{div}(\textbf{{grad}}^{\perp}(\psi)\,\,\omega)\,\textup{d}x+\mu\int_{\Omega}\phi_{2}\,\Delta\omega\,\textup{d}x.\end{aligned}\right$. (44)
@@
$\left\{\begin{aligned} \rho_{0}\int_{\Omega}\textbf{{grad}}(\phi_{1})\cdot\partial_{t}\textbf{{grad}}(\psi)\,\textup{d}x&={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\partial_{t}\,\bm{n}\cdot\textbf{{grad}}(\psi)\,\textup{d}s}+\int_{\Omega}\rho_{0}\,\omega\,\,\textbf{{grad}}(\phi_{1})\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}x\\ &\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\,\omega\,\bm{n}\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}s}-\mu\int_{\Omega}\Delta\phi_{1}\,\Delta\psi\,\textup{d}x\\ &\quad{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\mu\int_{\partial\Omega}\phi_{1}\,\bm{n}\cdot\textbf{{grad}}(\Delta\psi)\,\textup{d}s+\mu\int_{\partial\Omega}\Delta\psi\,\bm{n}\cdot\textbf{{grad}}(\phi_{1})\,\textup{d}s},\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\textbf{{grad}}^{\perp}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\,\psi\,\textup{d}x{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\,-\int_{\partial\Omega}\rho_{0}\,\phi_{2}\,\psi\,\bm{t}\cdot\textbf{{grad}}(\omega)\,\textup{d}s}\\ &\quad-\mu\int_{\Omega}\textbf{{grad}}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\textup{d}x{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}+\mu\int_{\partial\Omega}\phi_{2}\,\bm{n}\cdot\textbf{{grad}}(\omega)\,\textup{d}s}.\end{aligned}\right$. (45)
@@
$\left\{\begin{aligned} \rho_{0}\int_{\Omega}\textbf{{grad}}(\phi_{1})\cdot\partial_{t}\textbf{{grad}}(\psi)\,\textup{d}x&={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{0}\,\int_{\partial\Omega}\phi_{1}\,\partial_{t}\,u_{L}\,\textup{d}s}+\int_{\Omega}\rho_{0}\,\omega\,\,\textbf{{grad}}(\phi_{1})\cdot\textbf{{grad}}^{\perp}(\psi)\,\textup{d}x\\ &\quad+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\int_{\partial\Omega}\rho_{0}\,\phi_{1}\,\,\omega\,u_{D}^{1}\,\textup{d}s}-\mu\int_{\Omega}\Delta\phi_{1}\,\Delta\psi\,\textup{d}x\\ &\quad{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\mu\int_{\partial\Omega}\phi_{1}\,u_{D}^{2}\,\textup{d}s+\mu\int_{\partial\Omega}u_{D}^{3}\,\bm{n}\cdot\textbf{{grad}}(\phi_{1})\,\textup{d}s},\\ \rho_{0}\int_{\Omega}\phi_{2}\partial_{t}\omega\,\textup{d}x&=-\int_{\Omega}\rho_{0}\,\textbf{{grad}}^{\perp}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\,\psi\,\textup{d}x{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+\int_{\partial\Omega}\rho_{0}\,\phi_{2}\,\psi\,u_{D}^{4}\textup{d}s}\\ &\quad-\mu\int_{\Omega}\textbf{{grad}}(\phi_{2})\cdot\textbf{{grad}}(\omega)\,\textup{d}x{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}+\mu\int_{\partial\Omega}\phi_{2}\,u_{D}^{5}\,\textup{d}s}.\end{aligned}\right$. (46)
@@ 4.3.2 Finite Element approximation
Due to the dissipation term $\int_{\Omega}\Delta\psi\Delta\phi_{1}\,\textup{d}x$, $\psi$ has to be in $H^{2}$, hence let us choose an Argyris finite element family $(\phi^{1}_{i})_{i\in[1,N_{\psi}]}$. Given that Argyris elements are polynomials of degree 5 and the relation $\Delta\psi=\omega$, let us choose $H^{1}$ conforming $P^{3}$ Lagrange finite element family $(\phi^{2}_{i})_{i\in[1,N_{\omega}]}$ over $\Omega$. The discretized state variables are denoted by $\hat{\omega},\hat{\psi}$ and are defined as: $\hat{\psi}(t,x)=\sum_{i=1}^{N_{\psi}}\phi_{i}^{1}(x)\overline{\psi}_{i}(t)$, $\hat{\omega}(t,x)=\sum_{i=1}^{N_{\omega}}\phi^{2}_{i}(x)\overline{\omega}_{i}(t)$.
@@
$\left\{\begin{aligned} \rho_{0}\,\frac{\textup{d}}{\textup{d}t}\,\begin{bmatrix}\bm{K}&0\\ 0&\bm{M}\end{bmatrix}\begin{pmatrix}\overline{\psi}\\ \overline{\omega}\end{pmatrix}=&\begin{bmatrix}\begin{bmatrix}\bm{D}^{1}(\overline{\omega})&0\\ 0&-\bm{D}^{2}(\overline{\psi})\end{bmatrix}-\mu\begin{bmatrix}\bm{R}^{1}&0\\ 0&\bm{R}^{2}\end{bmatrix}\end{bmatrix}\begin{pmatrix}\overline{\psi}\\ \overline{\omega}\end{pmatrix}\\ &+\begin{bmatrix}\mu\,\bm{B}^{1}&\mu\,\bm{B}^{3}&\bm{B}^{3}\\ 0&0&0\end{bmatrix}\begin{pmatrix}u_{D}^{1}\\ u_{D}^{5}\\ \tilde{u}_{D}\end{pmatrix}+\mu\begin{bmatrix}0\\ \bm{B}^{5}\end{bmatrix}u_{D}^{5},\\ 0=&-\bm{B}^{1\top}\overline{\psi},\qquad\text{(No slip B.C.)}\\ 0=&-\bm{B}^{3\top}\overline{\psi},\qquad\text{(Impermeable B.C.)}\\ \bm{M}^{\partial}u_{D}^{1}=&\bm{B}^{5\top}\overline{\omega},\qquad\text{(Vorticity generation)}\end{aligned}\right$. (50)